OpenStudy (anonymous):

Olivia has $250 to buy presents for her 3 sisters and 2 nephews. She wants to spend $10 more on each sister than she spends on each nephew. Let s = the amount she spends on each sister, and n = the amount she spends on each nephew. How much (in dollars) does she spend on each sister? @texaschic101 @hartnn

hartnn (hartnn):

'spend $10 more on each sister than she spends on each nephew' ------> s = n+10

hartnn (hartnn):

Olivia has $250 to buy presents for her 3 sisters and 2 nephews 3 sisters ----3s 2 nephews----2n total = 250 3s+2n = 250

hartnn (hartnn):

plug in s= n+10 in this 2nd equation, you will only have an equation in 'n' then :)

OpenStudy (anonymous):

Oh, okay thanks! :)